The freezing point of water is 0.00°C at 1 atmosphere. How many grams of nickel(II) sulfate (154.8 g/mol), must be dissolved in 210.0 grams of water to reduce the freezing point by 0.400°C ? Refer to the table for the necessary boiling or freezing point constant. Solvent Formula Kb (°C/m) Kf (°C/m) Water H2O 0.512 1.86 Ethanol CH3CH2OH 1.22 1.99 Chloroform CHCl3 3.67 Benzene C6H6 2.53 5.12 Diethyl ether CH3CH2OCH2CH3 2.02 g nickel(II) sulfate.
Answer #1

NISO4(aq) ---------------> Ni^2+ (aq) + SO4^2- (aq)

Tf   = 0.40C

Kf     = 1.860C/m

i = 2

Tf    = i*Kf*m

   0.4    = 2*1.86*m

m = 0.4/(2*1.86) = 0.1075m

molality   = W*1000/G.M.Wt * weight of water

0.1075     = W*1000/(154.8*210)

W    = 0.1075*(154.8*210)/1000    = 3.5g

The weight of nickel(II) sulfate = 3.5g >>>>answer